OpenStudy (anonymous):

write the standard form of the equation of the parabola whose directrix is x=-1 and whose focus is at (5,-2)

OpenStudy (anonymous):

notice that your directrix is a vertical line so your equation will have this form: \[\huge 4p(x-h)=(y-k)^2 \] where (h,k) is the coordinates of your vertex and p is the distance from your vertex to your focus.....

OpenStudy (anonymous):

is the answer (y+2)^2=12(x+2)

OpenStudy (anonymous):

hint: the vertex is the midpoint between the focus and the directrix.... idk what the answer is until i work it out.... i thought you were...